OpenStudy (anonymous):

obtain the general solution of auxiliary equation y'' - 8y' + 17y =0 y(0)=4 y'(0)=-1

OpenStudy (perl):

did you find the characteristic solution

OpenStudy (perl):

r^2 -8r + 17=0 r = 4+i , 4 - i

OpenStudy (perl):

ok so far?

OpenStudy (perl):

the general solution is y = e^(4x) (c1*cos x + c2*sinx)

OpenStudy (perl):

here is a site that explains it http://www.sosmath.com/diffeq/second/constantcof/constantcof.html

OpenStudy (perl):

y(0)=4 y(0) = e^(4*0) ( c1 cos(0) + c2 sin(0)) = 1 ( c1 + 0 ) Therefore c1 = 4 now take the derivative (use product rule)